OpenStudy (anonymous):

solve by using the square root property. (5m-6)^2=7

OpenStudy (anonymous):

take square root of both sides so we get ((5m-6)^2)^1/2 = (7)^1/2 So we will get 5m - 6 = sqrt(7) so just solve for m

OpenStudy (anonymous):

try to move 6 to right hand side & also 5